# UFO: A Unified Approach to Fine-grained Visual Perception via Open-ended Language Interface

### v2

@@ 5.3 Fine-grained Instruction Tuning Results
$N^{2}$ 1 4 16 25
mAP 38.9 41.3 42.6 42.9
FPS 7.0 5.9 3.6 2.8
@@
Table 8: Ablation on COCO detection and instance segmentation.
5.4 Ablation Study
Number of Mask tokens. We ablate the number of mask tokens on the COCO instance segmentation task. As shown in Table 7, using multiple tokens significantly improves performance compared to a single token, but the gains plateau after 16. Considering the increased training and inference costs, we set $N^{2}=16$ by default for a balanced trade-off.
Adavanced training strategies. The sparsity of positive samples in multi-prediction tasks hampers effective learning. To mitigate this, we use two advanced strategies to increase the ratio of positive samples. First, copy-paste data augmentation [22], where objects from other images are pasted onto the target image. Second, we repeat ground truth $k$ times [26], defaulting to 3. As seen in Table 8, copy-paste boosts mAP by 1.1 for detection and 1.2 for instance segmentation, while repeating ground truth further boosts mAP by 1.6 and 2.2. This demonstrates that the sparsity of is a key bottleneck, and our performance can be effectively improved with these strategies. By default, we only use these strategies for COCO detection and instance segmentation.
6 Conclusion
In this paper, we present UFO, a unified approach for various fine-grained visual perception tasks with an open-ended language interface. We translate all perception targets into open-ended text sequences and introduce a novel embedding retrieval method for segmentation. Experiments show that our method can achieve excellent performance on MLLMs without requiring architecture modifications. Our unification fully aligns with vision-language tasks, providing a flexible, effective, and scalable solution to enhance the fine-grained perception capabilities of MLLMs, paving the way to build more general multimodal models.
Supplementary Material
Our supplementary material provides detailed information, including implementation details (§7), dataset descriptions (§8), and training specifics (§9). Inference speed and additional ablation studies are covered in §10 and §11, respectively. Experiments on more tasks and discussions with previous work are included in §12 and §13. Qualitative results from two training setups and visualizations of multiple mask tokens are presented in §14.
7 Implementation Details
Grid Generation. As mentioned in the multi-task template (Section 3.4), we divide multi-prediction tasks into sub-tasks, each corresponding to the nearest grid point. The number of grid points is roughly proportional to the image resolution, as shown in Table 11. Calculating the loss for all sub-tasks is computationally expensive. Therefore, we sample grid points and compute the loss on a subset, as presented in Table 11. When sampling, we prioritize positive samples. If the number of positive samples is less than the target size, we then sample from negative samples. During multi-task training on InternVL2-8B, we adjust the resolution based on the default resolution of the image tokenizer (using multiples of 448) and modify the grid point configurations accordingly. In fine-grained instruction tuning, we use the default resolution of the image tokenizer, such as $448^{2}$ for InternVL2 and $336^{2}$ for LLaVA-1.5. For the multi-prediction tasks of the two MLLMs, we use 100 grids and sample 40 of them.
Coordinate Discretization. We follow GiT to convert continuous coordinates into discrete numbers within the range [0, range]. Specifically, we define the range as twice the image resolution. For example, for a 448 $\times$ 448 image, we convert coordinate values into integers within [0, 896].
Data Augmentation. For object detection, instance and sematic segmentation, we use RandomFlip and RandomResizedCrop. We also use CopyPaste for object detection and instance segmentation. For other tasks, we simply resize the images to the required resolution.
Indicator function. In segmentation, we use sigmoid as the indicator and then get a mask with a 0.5 threshold.
8 Dataset Details
Fine-grained Instruction Tuning Datasets As shown in Table 9, our Fine-grained Instruction Tuning datasets build upon GiT [61]’s universal datasets with the following modifications: (1) Replacing captioning datasets with VQA datasets to maintain the multimodal conversation capabilities of MLLMs. (2) Excluding datasets with smaller data volumes, such as Pascal VOC and BDD100K. (3) Following LISA [33], adding PACO-LVIS and Pascal Part to enhance segmentation of object parts.
9 Training Details
Multi-Task Training Details. Multi-task training setting is in Table 10. For single-task training, we simply reduce the training iterations to 120k. For multi-task training on InternVL2-8B, we keep all parameters trainable and reduce the iterations to 400k because of its faster convergence.
Progressive High-resolution Training of MLLMs. In multi-task training, object detection, instance segmentation, and semantic segmentation require predicting a large number of targets, including many small objects, making performance highly sensitive to resolution. For the relatively small UFO-ViT, we train directly using high resolution. However, for multi-task training on MLLMs, high resolution significantly increases training costs. Therefore, we adopt a progressive high-resolution training strategy: first training at a resolution of $448^{2}$ for 300k iterations, then at $896^{2}$ for 60k iterations, and finally at $1344^{2}$ for 40k iterations. We utilize InternVL2-8B’s dynamic resolution to support high-resolution inputs. As shown in Table 12, increasing the resolution leads to substantial improvements in detection and segmentation performance, even with fewer iterations.
Fine-grained Instruction Tuning Details. Training setting is in Table 10. The training data for the MLLM includes six tasks, each containing multiple datasets. We first sample evenly from the six tasks, with a sampling ratio of 1/6 for each task. Then, within each task, we sample evenly from the corresponding datasets. For example, for object detection, the sampling ratio for each dataset is 1/30. When fine-tuning on a specific dataset, we maintain the same training setup but train for only 20k iterations.
Task Sources Size
VQA LLaVA-v1.5- mix665k [39] 0.7m
 RefCOCO [75], RefCOCO+ [75],
REC & RES  RefCOCOg [75], RefCLEF [28] 0.8m
Objects365 [56], OpenImages [32],
Object Detection  COCO [37], LVIS [23], nuImages [4] 3.8m
OpenImages [32], LVIS [23],  COCO [37],
Instance Seg PACO-LVIS [50], PASCAL-Part [9], nuImages [4], 1.4m
COCOStuff [3],Mapillary [44]
Semantic Seg nuImages [4],, ADE20K [80] 0.3m
Table 9: Fine-grained instruction tuning datasets.
config ViT MLLM MLLM-instruct
optimizer AdamW AdamW AdamW
learning rate 2e-4 2e-4 2e-4
weight decay 0.05 0.01 0.01
layer-wise lr decay 0.85 0.85 -
schedule cosine cosine cosine
gradient norm clip 0.1 1.0 1.0
warmup iters 1k 1k 1k
training iters 640k 400k 120k
batch size 24 24 32
gradient accumulation - - 16
LoRA rank - - 8
LoRA alpha - - 16
LoRA dropout - - 0.05
LoRA modules - - LLMs
drop path 0.1(B), 0.4(L,H) - -
precision FP16 BF16 BF16
GPUS 24 $\times$ V100 8$\times$ A100 8$\times$ A100
Table 10: Multi-task training and instruction tuning settings.
Task Resolution Grid Sample Grid Speed
Object Detection $1120^{2}$ 625 250 4.1
Instance Seg $1120^{2}$ 625 250 3.6
Semantic Seg $672^{2}$ 225 90 4.8
Image Captioning $224^{2}$ 0 0 7.7
REC $224^{2}$ 0 0 9.1
Resolution Iters Detection Ins Seg Sem Seg
$448^{2}$ 300k 44.3 37.4 53.9
$896^{2}$ 60k 51.7 44.1 54.6
$1344^{2}$ 40k 52.3 45.8 -
Table 12: Iterations and performance in progressive high-resolution training on UFO-InternVL2-8B.
Model   REC  RES  ReasonSeg
UFO-InternVL2-8B 1.10 0.58 0.57
UFO-LLaVA-1.5-7B 0.78 0.67 0.65
10 Inference Speed
Table 11 shows the speed of UFO-ViT-B. By using parallel decoding for multi-prediction tasks, we achieve inference speeds comparable to single-prediction tasks, despite higher resolutions ($1120^{2}$ vs. $224^{2}$) and more predictions. Table 13 shows the speed on MLLMs. Our LLaVA-1.5 variant is slower than InternVL2 on REC because its tokenizer converts textual numbers into longer token sequences. In embedding retrieval, the extra scaled-dot product operation only costs a negligible 0.17 ms for InternVL2-8B on an A100.
Beam Detection Instance Seg Captioning
Number mAP mAP BLEU-4 CIDEr
1 45.6 40.9 33.0 108.5
2 47.4 42.1 34.2 111.1
3 47.8 42.6 34.0 110.8
5 47.9 42.6 33.9 111.0
[…]
CE:Focal 1:1 1:3 3:1
Ins Seg 43.5 43.7 43.4
Caption 35.3 34.9 35.3
@@ 10 Inference Speed
box 90.7 1.10
mask2box 89.5 0.57
Table 16: REC performance.
RefCOCO val RefCOCO val ReasonSeg
REC RES VQA Ins Seg Sem Seg Det REC RES test
88.2 - -
88.6 73.5 47.7
89.2 74.4 50.2
89.6 76.7 54.9
90.7 77.3 55.4
Table 17: Ablation of MLLM training data on UFO-InternVL2-8B.
11 More ablation studies
@@
Loss weights. In Table 16, we ablate loss weights on UFO-ViT-B. We jointly training both instance segmentation and image captioning tasks. Although increasing the focal loss weight is slightly better for segmentation, it leads to a drop in captioning. Since our goal is better overall multi-task performance, we set all weights to 1 to avoid task competition.
Data contributions. In Table 17, we illustrate the impact of different task data on performance. We observe that adding VQA data significantly enhances performance on ReasonSeg, likely because it helps maintain the language reasoning abilities of MLLMs. The inclusion of instance and semantic segmentation data benefits both RES and ReasonSeg by providing dense mask annotations. Additionally, incorporating detection data leads to noticeable improvements in the REC task, as it offers dense bounding box annotations and potentially exposes the model to a broader range of concepts.
Comparisons with baseline MLLMs. In Table 18, we provide comparisons between UFO and the baseline MLLM. Firstly, UFO significantly expands the task range, enabling the model to handle all types of segmentation. Secondly, although InternVL2-8B can support REC and perform detection by breaking it into multiple single-category prediction tasks, its performance is markedly inferior to us, especially in detection. This is primarily because InternVL2-8B cannot predict multiple boxes for a single category nor model the relationships among multiple categories, leading to insufficient and contradictory predictions. In contrast, UFO effectively supports multi-prediction tasks through local prompts, allowing it to accommodate any prediction number.
Model Det Ins Seg Sem Seg REC RES
InternVL2-8B [9] 10.5 - - 87.1 -
UFO-InternVL2-8B 52.3 45.8 54.6 90.7 77.3
@@
12 Extended Experiments
Retinal Vessel Segmentation. Our embedding retrieval method offers superior expressive capability compared to polygons, particularly in highly complex and detailed masks, which require a large number of vertices when using polygons. To further illustrate this, we fine-tune our model on the retinal vessel segmentation, where the vessels possess very irregular and narrow shapes, which are hard to represent as polygons. We follow the few-shot settings in GiT [61], fine-tuning both UFO-ViT-H and UFO-InternVL2-8B on the DRIVE [58] training set for only 100 steps. As shown in Figure 5, UFO accurately segments the retinal vessels. In performance, UFO-ViT-H achieves 77.4 Dice, outperforming GiT-H with 57.9 Dice. UFO-InternVL2-8B also achieves a competitive 76.3 Dice. This result validates the effectiveness of our segmentation method on extremely fine-grained structures, enabling support for more general segmentation.
Figure 5: Visualizations of retinal vessel segmentation.
Depth Estimation. Thanks to the flexibility of our method, we can easily extend it to depth estimation similar to segmentation. We can apply a sigmoid to the dot product result to interpret it as relative depth $\boldsymbol{r}$, which can be then mapped into absolute depth. For depth $\hat{\mathbf{D}}$ within [$\mathbf{D}_{min}$, $\mathbf{D}_{max}$], we can predict it as follows:
$\boldsymbol{r}=\sigma(\frac{\mathbf{e_{d}}\,\mathbf{h_{v}}^{\top}}{\sqrt{d}}),\quad\hat{\mathbf{D}}=r\cdot\mathbf{D}_{max}+(1-r)\cdot\mathbf{D}_{min}\vspace{-3pt}$ (6)
$\mathbf{e_{d}}$ is the embedding of <DEPTH> token. As shown in Table 19, we can achieve competitive performance.
@@ 12 Extended Experiments
UFO-InternVL2-8B 0.320 0.928 0.091 0.039
Table 19: Depth estimation performance on NYUv2 Depth [57].
13 Discussions
Comparisons with GiT. GiT [61] also aims to build a generalist model for fine-grained perception tasks. Compared with GiT, we provides six key improvements: 1) Segmentation by embedding retrieval, a simple yet intuitive way to support segmentation by language interface. GiT uses polygons and textual classes, leading to information loss or lengthy sequences. In contrast, UFO can accurately segment using only 16 mask tokens. 2) Alignment with the open-ended language interface: unlike GiT, which requires separate vocabularies and fixed output lengths per task, UFO uses shared vocabulary and outputs arbitrary-length sequences. Tables 6 and  18 demonstrate that open-ended detection is challenging due to severe class imbalance. We address this issue with a text-aligned beam search and achieve enhanced performance. 3) Scalability to MLLMs: while GiT only experiments on relatively small ViTs, UFO can easily scale to larger MLLMs thanks to the aligned language interface. 4) Exploring the image representation capabilities of the language interface. GiT is a purely text-based method, while UFO can effectively extract mask information from image features. 5) Better task universality: GiT uses different methods for instance and semantic segmentation (polygon and textual class), while we adopt a unified approach for both tasks because UFO can support masks with any shape. As UFO is aligned with vision-language tasks, we can effortlessly combine VQA reasoning and segmentation, enabling ReasonSeg. 6) Significant performance improvements. In Table 2, UFO-ViT-H outperforms GiT-H by 1.2 mAP and 12.3 mAP on COCO detection and instance segmentation, and 3.3 mIoU on ADE20K semantic segmentaion.
14 Visualization
@@
Instruction Tuning Results. We present qualitative results of UFO-InternVL2-8B in Figue 8. Leveraging the language capabilities of MLLMs, the model can accurately locate and segment based on both simple phrases and complex queries.
@@
Failure cases. We visualize the failure cases of UFO-InternVL2-8B on REC, RES and ReasonSeg in Figure 6. In the first example, the model misunderstands “fourth” and detects the third guy from the right, showing its inaccuracy in counting. In the second example, the model only marks visible table parts (blocked by food), while the human guesses hidden areas. This is mainly because the semantics of image features on occluded areas are different. However, whether to include occluded areas is also ambiguous. In the third example, the model fails to identify the answer as “river” due to limited knowledge, leading to segmentation errors.
Figure 6: Failure case visualizations of UFO-InternVL2-8B on REC, RES and ReasonSeg respectively.
@@
Figure 9: Visualization of multiple mask tokens. We illustrate with four mask tokens (with $N$=2). Employing multiple mask tokens allows for capturing finer details, such as the horse leg and the dog tail, resulting in more precise and refined masks.

### v3

@@ Appendix D Training Details
Table 13: Comparison of trainable parameters when applying LoRA.
Method Base (M)LLM LoRA Rank LoRA Parameters Text Embedding
Cores-7B [3] LLaVA-7B [43] 8 20M 262M
GLaMM-7B [57] Vicuna-7B [14] 8 20M 262M
SAM4MLLM-8B [11] Qwen-VL-7B [1] 256 693M 1.24B
UFO-LLaVA-1.5-7B LLaVA-1.5-7B [43] 8 20M 262M
UFO-InternVL2.5-8B InternVL2.5-8B [12] 8 19M 758M
Appendix E Inference Speed
Table 9 shows the speed of UFO-ViT-B. By using parallel decoding for multi-prediction tasks, we achieve inference speeds comparable to single-prediction tasks, despite higher resolutions ($1120^{2}$ vs. $224^{2}$) and more predictions. Table 11 shows the speed on MLLMs. Our LLaVA-1.5 variant is slower than InternVL2.5 on REC because its tokenizer converts textual numbers into longer token sequences. In embedding retrieval, the extra scaled-dot product operation only costs a negligible 0.17 ms for InternVL2.5-8B on an A100.
@@ Appendix E Inference Speed
Beam Detection Instance Seg Captioning
Number mAP mAP BLEU-4 CIDEr
1 45.6 40.9 33.0 108.5
2 47.4 42.1 34.2 111.1
3 47.8 42.6 34.0 110.8
5 47.9 42.6 33.9 111.0
Table 15: Ablation on COCO detection and instance segmentation.
@@
CE:Focal 1:1 1:3 1:5 3:1
Instance Seg 43.5 43.7 43.8 43.4
Captioning 35.3 34.9 34.8 35.3
Table 17: Ablation of direct box prediction.
@@
box 91.8 1.10
mask2box 90.5 0.57
Appendix F More ablation studies
@@
Loss weights. In Table 17, we ablate loss weights on UFO-ViT-B. We jointly train both instance segmentation and image captioning tasks. Although increasing the focal loss weight is slightly better for segmentation, it leads to a drop in captioning. Since our goal is better overall multi-task performance, we set all weights to 1 to avoid adding task bias.
Advanced training strategies. The sparsity of positive samples in multi-prediction tasks hampers effective learning. To mitigate this, we use two advanced strategies to increase the ratio of positive samples. First, copy-paste data augmentation [23], where objects from other images are pasted onto the target image. Second, we repeat ground truth $k$ times [29], defaulting to 3. As seen in Table 15, copy-paste boosts mAP by 1.1 for detection and 1.2 for instance segmentation, while repeating ground truth further boosts mAP by 1.6 and 2.2. This demonstrates that the sparsity of is a key bottleneck, and our performance can be effectively improved with these strategies. By default, we only use these strategies for COCO detection and instance segmentation.
Comparisons with baseline MLLMs. In Table 18, we provide comparisons between UFO and the baseline MLLM. Firstly, UFO significantly expands the task range, enabling the model to handle all types of segmentation. Secondly, although InternVL2.5-8B can support REC and perform detection by breaking it into multiple single-category prediction tasks, its performance is markedly inferior to ours, especially in detection. This is primarily because InternVL2.5-8B cannot predict multiple boxes for a single category nor model the relationships among multiple categories, leading to insufficient and contradictory predictions. In contrast, UFO effectively supports multi-prediction tasks through local prompts, allowing it to accommodate any prediction number.
@@
Model Detection Instance Seg Semantic Seg REC RES MMVP HallBench
InternVL2.5-8B [12] 12.5 - - 90.3 - 76.3 50.1
UFO-InternVL2.5-8B 52.3 45.8 54.6 93.1 81.0 76.3 50.7
@@
Appendix G Extended Experiments
Retinal Vessel Segmentation. Our embedding retrieval method offers superior expressive capability compared to polygons, particularly in highly complex and detailed masks, which require a large number of vertices when using polygons. To further illustrate this, we fine-tune our model on the retinal vessel segmentation, where the vessels possess very irregular and narrow shapes, which are hard to represent as polygons. We follow the few-shot settings in GiT [69], fine-tuning both UFO-ViT-H and UFO-InternVL2.5-8B on the DRIVE [64] training set for only 100 steps. In performance, UFO-ViT-H achieves 77.4 F1 score, outperforming GiT-H with 57.9 score. UFO-InternVL2.5-8B also achieves a competitive 78.1 F1 score. After fine-tuning with the entire training set, our performance can surpass strong specialized models such as UNet++ [91] and ConvMixer [50]. As shown in Figure 5, UFO accurately segments the retinal vessels. This result validates the effectiveness of our method on extremely fine-grained structures, enabling support for more general segmentation.
Table 19: DRIVE [64] Segmentation.
Methods 5-shot Full fine-tuning
UNet++ [91] - 79.6
ConvMixer [50] - 82.2
GiT-H [69] 57.9 -
UFO-ViT-H 77.4 82.0
UFO-InternVL2.5-8B 78.1 82.4
Table 20: Depth estimation on NYUv2 Depth [63].
@@ Appendix G Extended Experiments
UFO-InternVL2.5-8B 0.305 0.936 0.087 0.035
Figure 5: Visualizations of retinal vessel segmentation.
Depth Estimation. Thanks to the flexibility of our method, we can easily extend it to depth estimation similar to segmentation. We can apply a sigmoid to the dot product result to interpret it as relative depth $\boldsymbol{r}$, which can be then mapped into absolute depth. For depth $\hat{\mathbf{D}}$ within $\mathbf{D}_{min}$, $\mathbf{D}_{max}$], we can predict it as follows:
$\boldsymbol{r}=\sigma(\frac{\mathbf{e_{d}}\,\mathbf{h_{v}}^{\top}}{\sqrt{d}}),\quad\hat{\mathbf{D}}=r\cdot\mathbf{D}_{max}+(1-r)\cdot\mathbf{D}_{min}\vskip-3.0pt$ (8)
$\mathbf{e_{d}}$ is the embedding of <DEPTH> token. As shown in Table 20, we can achieve competitive results.
Surface Normal Prediction. Similar to depth estimation, we can extend UFO to surface normal prediction. We introduce three task-specific tokens for normal vectors: <NORMAL_X>, <NORMAL_Y>, and <NORMAL_Z>. Given the image feature $\mathbf{h}_{v}$ and the three directional embeddings $\mathbf{e}_{x}$, $\mathbf{e}_{y}$, and $\mathbf{e}_{z}$, the normal components (e.g., $\hat{n}_{x}$) are computed as follows:
$\hat{n}_{x}=\sigma\left(\frac{\mathbf{e}_{x}^{\top}\mathbf{h}_{v}}{\sqrt{d}}\right)$
where $\sigma(\cdot)$ denotes the sigmoid function and $d$ is the feature vector dimension. Finally, the predicted values are normalized to obtain a unit surface normal vector:
$\mathbf{\hat{n}}=\frac{(\hat{n}_{x},\,\hat{n}_{y},\,\hat{n}_{z})}{\left\|(\hat{n}_{x},\,\hat{n}_{y},\,\hat{n}_{z})\right\|}$
We conduct the evaluation on NYU v2 Normal Benchmark, and the performance is shown in the Table 21. It can be seen that our method achieves comparable performance to specialized models.
Table 21: Surface normal prediction on NYUv2 Depth [63]
Method Mean Median $11.25^{\circ}$ $22.5^{\circ}$ $30^{\circ}$
GeoNet++ [55] 18.5 11.2 9.502 0.732 0.907
Marigold [32] 18.8 - 0.559 - -
GeoWizard [22] 17.0 - 0.565 - -
UFO-InternVL2.5-8B 17.8 10.4 0.543 0.733 0.800
Appendix H Discussions
Comparisons with GiT. GiT [69] also aims to build a generalist model for fine-grained perception tasks. Compared with GiT, we provide six key improvements: 1) Segmentation by embedding retrieval, a simple yet intuitive way to support segmentation by language interface. GiT uses polygons and textual classes, leading to information loss or lengthy sequences. In contrast, UFO can accurately segment using only 16 mask tokens. 2) Alignment with the open-ended language interface: unlike GiT, which requires separate vocabularies and fixed output lengths per task, UFO uses shared vocabulary and outputs arbitrary-length sequences. Tables 7 and  18 demonstrate that open-ended detection is challenging due to severe class imbalance. We address this issue with a text-aligned beam search and achieve enhanced performance. 3) Scalability to MLLMs: while GiT only experiments on relatively small ViTs, UFO can easily scale to larger MLLMs thanks to the aligned language interface. 4) Exploring the image representation capabilities of the language interface. GiT is a purely text-based method, while UFO can effectively extract mask information from image features. 5) Better task universality: GiT uses different methods for instance and semantic segmentation (polygon and textual class), while we adopt a unified approach for both tasks because UFO can support masks with any shape. As UFO is aligned with vision-language tasks, we can effortlessly combine VQA reasoning and segmentation, enabling ReasonSeg. 6) Significant performance improvements. In Table 2, UFO-ViT-H outperforms GiT-H by 1.2 mAP and 12.3 mAP on COCO detection and instance segmentation, and 3.3 mIoU on ADE20K semantic segmentation.
Appendix I Visualization
@@
Instruction Tuning Results. We present qualitative results of UFO-InternVL2.5-8B in Figure 7. Leveraging the language capabilities of MLLMs, the model can accurately locate and segment based on both simple phrases and complex queries.
@@
Failure cases. We visualize the failure cases of UFO-InternVL2.5-8B on REC, RES and ReasonSeg in Figure 10. In the first image, the arm that needs to be localized is very blurry and only appears in a small portion at the edge of the image, indicating that the model has deficiencies in localizing objects with low visibility. In the second image, multiple zebras overlap and their patterns are very similar, resulting in incorrect segmentation locations, demonstrating the model’s shortcomings in segmenting clustered objects. In the third example, the model fails to identify the answer as “river" due to limited knowledge, leading to segmentation errors.
Appendix J Broader Impacts
Our approach achieves a compact model design by removing task-specific decoders and integrating diverse fine-grained perception tasks into a unified architecture. This structural efficiency reduces computational demands during training, leading to a decrease in carbon footprint. We do not identify negative social impacts currently.
@@
Figure 8: Visualization of multiple mask tokens. We illustrate with four mask tokens (with NN=2). Employing multiple mask tokens allows for capturing finer details, such as the horse leg and the dog tail, resulting in more precise and refined masks.
@@
Figure 10: Failure case visualizations of UFO-InternVL2.5-8B on REC, RES and ReasonSeg.